# Beyond the Hype: A dispassionate look at vision-language models in medical scenario

### v1

@@ VI Discussion
Overall Performance. The overall performance of LVLMs remains stable on both CT and MRI datasets, with GPT, Gemini, Huatuo (7b and 34b), and Claude achieving the top-5 performance of all comparisons. Specifically, GPT-4o achieved the highest capabilities on both CT and MRI datasets. Huatuo-34 also shows reliable performance in OEQ, making it a strong competitor across different modalities. While models like Gemini and Claude show more variability in their rankings, their presence in the top five for both datasets suggests their high effectiveness.
The result also suggests that commercial LVLMs continue to outperform their open-source counterparts. Although LLaVA-Med is purportedly better suited for the medical domain, its performance still falls short of generalized open-source LVLMs. Notably, Huatuo-34b demonstrates comparable performance, approaching that of GPT-4o on CT data, despite its smaller model scale. This highlights the potential for medical-specific LVLMs to surpass commercial models as they scale up.
Is LVLM sensitive to CT windowing? The performance of LVLMs on CT data was widely assessed in existing benchmark studies. Unfortunately, none of them considers the windowing settings while barely collecting 2D CT images (those have been preprocessed) or simply applying min-max normalization. CT windowing is widely used to adjust the contrast and brightness of the images, enhancing the visibility of different tissues and structures within the body. By adjusting the windowing parameters (window level WL and window width WW), radiologists can better differentiate between normal and abnormal tissues, resulting in better diagnosis/screening efficiency.
@@
Table IV presents the windowing effects on LVLMs, with the values representing the residual in open-ended(hallucination) scores between results with CT windowing and those without CT windowing. Interestingly, the differences between w/o windowing and with windowing were almost consistent. For instance, applying windowing significantly (p<0.05) helped LVLMs to improve anatomic understanding (Huatuo-7b, Huatuo-34b, InternVL, MiniCPM, GPT-4o, Claude) and spatial reasoning (LLaVA-Med, LLaVA, MiniCPM, Gemini, Claude), while it might have opposite effects for certain models. In particular, the anatomic understanding and physiological knowledge scores of BLIP2 were notably decreased when applying CT windowing, with -0.13 and -0.49 RA scores, respectively. This suggests that BLIP2 may be trained on pure natural data without any medical samples.
@@
As Table V shows, most LVLMs (LLaVA-Med, LLaVA, MiniCPM, BLIP2, and GPT-4o) benefited from prompt CoT in open-ended questions, with higher RA scores (around 0.2 average gain) and lower hallucination frequencies. On the contrary, more hallucinations were observed in certain LVLMs (with 4.4% of Claude-Sonnet and 0.41% of Gemini-Flash), indicating that CoT may also mislead LVLMs to misjudgments. For closed-ended questions, in addition to the InternVL and Huatuo variants, the MCA scores of all other LVLMs significantly (p<0.05) improved after introducing the chain of thought prompts.
@@
Are LVLMs robust to unharmonised data? Table III demonstrates the performance of LVLMs on unharmonised data. To further explore the effect of different variants, we compared the LVLMs’ performance (on unharmonised data, e.g., noise, blur, and contrast adjustments) against those on the raw data. Interestingly, we found that the tolerances of different LVLMs for unharmonised data were various. For instance, while Huatuo (both 7B and 34B) demonstrated reduced accuracy on augmented data, it experienced a decline in the rate of hallucinations. On the contrary, although Gemini achieved slightly higher accuracy scores in close-ended questions, it dramatically encountered hallucination issues (from an average hallucination of 0.7% to 24.28%). Notably, InternVL demonstrated the most stable results, with the least performance shift observed on the unharmonised data, indicating a strong robustness to data variation.
@@
This suggests that future research must address these gaps by systematically considering input data characteristics, therefore harmonising the input data to the same data space. This involves developing harmonisation foundation models that resolve the heterogeneity of medical imaging practices for multi-centre datasets. By doing so, LVLMs can not only handle in-domain queries but also achieve reliable and effective results in diverse clinical settings.
@@
Can LVLMs distinguish synthetic data from the real ones? To investigate whether LVLMs could distinguish synthesised data, we calculated the synthetic-detection rates of LVLMs. Our finding (Table VI) suggests that LVLMs struggle with synthetic data, even when the synthetic one does not follow the human physiological structures (Fig. 4). This is particularly true when the synthetic data is generated using advanced techniques that replicate the intricate patterns and textures of real data.
@@
Specifically, all LVLMs exhibited low synthetic-detection rates, ranging from 0% to 35.2%. However, it is worth noting that some models adopt a more sophisticated approach to handle this uncertainty. For example, GPT-4o stands out by refusing to answer nearly all queries related to synthetic data identification. This behaviour suggests a better safety mechanism, where the model opts for caution rather than providing potentially inaccurate information. Such a refusal to respond may be interpreted as an acknowledgement of the model’s limitations, allowing it to avoid the risk of hallucination or misclassification. Moreover, all LVLMs showed weak capabilities in distinguishing nonhuman data, with RA scores of 1.03 (BLIP2) to 2.27 (LLaVA-Med).
@@ VII Conclusion
  • LVLMs are not robust to unharmonised data, even certain models have lower hallucinations on the low-quality data.

### v2

@@ VI Discussion
Overall Performance. The overall performance of LVLMs remains stable on both CT and MRI datasets, with GPT, Gemini, Huatuo (7b and 34b), and Claude achieving the top-5 performance of all comparisons. Specifically, GPT-4o achieved the highest capabilities on both CT and MRI datasets. Huatuo-34 also shows reliable performance in OEQ, making it a strong competitor across different modalities. While models like Gemini and Claude show more variability in their rankings, their presence in the top five for both datasets suggests their high effectiveness. In particular, the most powerful LVLM (GPT-4o and Gemini) could only achieve comparable results in anatomic understanding and quantitative reasoning, while it underperformed in interpreting linguistic and visual instructions (multimodal comprehension). Interestingly, all models struggled with physiological knowledge, likely due to their limited ability to comprehend visual instructions. Meanwhile, the results indicate that small models (e.g., 7b scales) are prone to hallucination issues. These findings highlight a significant gap in the applicability of current LVLMs to clinical practice, particularly in scenarios requiring the integration of visual and linguistic instructions.
Regarding the trade-offs between model scale and domain-specific knowledge, our findings suggest that, while large closed-source models (e.g., GPT-4) benefit from vast parameter counts and broad general-domain exposure, domain-specific adaptations can help smaller or open-source models (like Huatuo-34b) close the performance gap. However, this relies on the quality and quantity of domain-specific data and may not apply in all cases. For instance, the comparison between LLaVA and its medical adaptation LLaVA -Med shows that even a model originally trained on natural data can achieve better results compared with the one receiving specialized medical data and instruction tuning. This observation indicates that domain-relevant data plays a crucial role—potentially as important as (or even more critical than) raw model scale—in elevating performance on medical imaging tasks. The result also suggests that commercial LVLMs continue to outperform their open-source counterparts. Although LLaVA-Med is purportedly better suited for the medical domain, its performance still falls short of generalized open-source LVLMs. Notably, Huatuo-34b demonstrates comparable performance, approaching that of GPT-4o on CT data, despite its smaller model scale. This highlights the potential for medical-specific LVLMs to surpass commercial models as they scale up.
Is LVLM sensitive to CT windowing? The performance of LVLMs on CT data was widely assessed in existing benchmark studies. Unfortunately, none of them considers the windowing settings while barely collecting 2D CT images (those have been preprocessed) or simply applying min-max normalization. CT windowing is a critical step in clinical practice for optimizing the display of specific tissues. For instance, lung windows highlight pulmonary parenchyma details, while mediastinal windows emphasize soft-tissue contrast. By adjusting the windowing parameters (window level WL and window width WW), radiologists can better differentiate between normal and abnormal tissues, resulting in better diagnosis/screening efficiency.
@@
Table IV presents the windowing effects on LVLMs, with the values representing the residual in open-ended(hallucination) scores between results with CT windowing and those without CT windowing. Interestingly, the trends between w/o windowing and with windowing were almost consistent. For instance, applying windowing significantly (p<0.05) helped LVLMs to improve anatomic understanding (Huatuo-7b, Huatuo-34b, InternVL, MiniCPM, GPT-4o, Claude) and spatial reasoning (LLaVA-Med, LLaVA, MiniCPM, Gemini, Claude), while it might have opposite effects for certain models. In particular, the anatomic understanding and physiological knowledge scores of BLIP2 were notably decreased when applying CT windowing, with -0.13 and -0.49 RA scores, respectively. This suggests that BLIP2 may be trained on pure natural data without any medical samples.
@@
As Table V shows, most LVLMs (LLaVA-Med, LLaVA, MiniCPM, BLIP2, and GPT-4o) benefited from prompt CoT in open-ended questions, with higher RA scores (around 0.2 average gain) and lower hallucination frequencies. On the contrary, more hallucinations were observed in certain LVLMs (with 4.4% of Claude-Sonnet and 0.41% of Gemini-Flash), indicating that CoT may also mislead LVLMs to misjudgments. Our findings indicate that Chain-of-Thought (CoT) prompting has a varied impact on different reasoning tasks, with significant improvements in certain areas while showing limited effects in others (details in the supplementary materials). Specifically, multimodal comprehension exhibited the most notable enhancement across all LVLMs, with an average increase of 0.40 in RA scores and a 5.59% reduction in hallucinations. Similarly, spatial reasoning and quantitative reasoning benefited from CoT prompting, with observed RA score gains of 0.28 and 0.29, respectively, and corresponding reductions in hallucinations by 5.1% and 4.29%. Interestingly, however, CoT did not lead to significant improvements in the models’ physiological knowledge, and the results for anatomic understanding were inconsistent across different models. One possibility is that CoT can occasionally mislead models by reinforcing incorrect intermediate reasoning steps, particularly when the underlying knowledge is uncertain or complex. For closed-ended questions, in addition to the InternVL and Huatuo variants, the MCA scores of all other LVLMs significantly (p<0.05) improved after introducing the chain of thought prompts.
While CoT can enhance reasoning capabilities in certain contexts, its effectiveness appears to vary widely, mainly due to the following reasons 1) The model scale limits the ability to fully exploit CoT reasoning. Small models face parameter capacity limitations, struggling to maintain long-range dependencies and intermediate reasoning steps due to restricted working memory and shallow attention layers. In contrast, larger, often closed-source models (e.g., 100B+ parameters) tend to benefit more substantially from CoT [27]. 2) For essential medical questions (e.g., basic factual queries), applying a chain of thought may not yield significant improvement for some approaches (e.g., Huatuo) since the model’s performance partially depends on knowledge gained during pretraining, rather than elaborate reasoning steps.
These findings highlight the need for a more adaptive CoT approach that tailors prompting strategies based on the complexity and nature of the reasoning task. For instance, leveraging dynamic prompt tuning—where CoT depth and structure are adjusted based on the model’s initial response confidence—could help mitigate potential pitfalls. Future work should explore strategies for optimizing CoT prompting by systematically analyzing failure cases and refining prompt structures to better align with different reasoning modalities. Understanding when and why CoT is beneficial—or detrimental—will be key to maximizing its potential across diverse clinical and multimodal AI applications.
@@
Are LVLMs robust to unharmonized data? Table III demonstrates the performance of LVLMs on unharmonized data. To further explore the effect of different variants, we compared the LVLMs’ performance (on unharmonized data, e.g., noise, blur, and contrast adjustments) against those on the raw data. Interestingly, we found that the tolerances of different LVLMs for unharmonized data were various. For instance, while Huatuo (both 7B and 34B) demonstrated reduced accuracy on augmented data, it experienced a decline in the rate of hallucinations. On the contrary, although Gemini achieved slightly higher accuracy scores in close-ended questions, it dramatically encountered hallucination issues (from an average hallucination of 0.7% to 24.28%). Notably, InternVL demonstrated the most stable results, with the least performance shift observed on the unharmonized data, indicating a strong robustness to data variation.
To further explore data harmonization effects, we applied histogram equalization as a preprocessing step (shown in Table III). Interestingly, most LVLMs showed considerable improvements in closed-ended question (CEQ) performance under this setting. Notably, Gemini-Flash reached a 0.784 MCA score—an increase of approximately 10%—making it the top-performing model for CEQs. However, the same preprocessing appeared to hinder open-ended question performance for all but BLIP2. This discrepancy may arise from BLIP2’s training procedure, which potentially included histogram-equalized images, thus enabling the model to adapt more readily. Overall, these results highlight that while histogram equalization can enhance certain tasks, its efficacy may vary depending on the model’s training protocol and the nature of the queries.
@@
This suggests that future research must address these gaps by systematically considering input data characteristics, therefore harmonizing the input data to the same data space. This involves developing harmonization foundation models that resolve the heterogeneity of medical imaging practices for multi-center datasets. By doing so, LVLMs can not only handle in-domain queries but also achieve reliable and effective results in diverse clinical settings.
@@
Can LVLMs distinguish synthetic data from the real ones? To investigate whether LVLMs could distinguish synthesized data, we calculated the synthetic-detection rates of LVLMs. Our finding (Table VI) suggests that LVLMs struggle with synthetic data, even when the synthetic one does not follow the human physiological structures (Fig. 4). This is particularly true when the synthetic data is generated using advanced techniques that replicate the intricate patterns and textures of real data.
@@
Specifically, all LVLMs exhibited low synthetic-detection rates, ranging from 0% to 35.2%. However, it is worth noting that some models adopt a more sophisticated approach to handle this uncertainty. For example, GPT-4o stands out by refusing to answer nearly all queries related to synthetic data identification. This behavior suggests a better safety mechanism, where the model opts for caution rather than providing potentially inaccurate information. Such a refusal to respond may be interpreted as acknowledging the model’s limitations, allowing it to avoid the risk of hallucination or misclassification. Moreover, all LVLMs showed weak capabilities in distinguishing nonhuman data, with RA scores of 1.03 (BLIP2) to 2.27 (LLaVA-Med).
Our findings suggest that current LVLMs struggle to reliably distinguish synthetic images, raising important questions about how to maintain data integrity in a clinical setting. The detection of synthetic data may be addressed in the following ways: 1) Assessing data distribution. Normally, synthetic data lacks the natural variability and noise present in real-world data, this includes: uniform distributions where real data would have peaks and valleys, unusual consistency in patterns, a lack of outliers, and easily identifiable repeating sequences; 2) Watermark detection. Using invisible, machine-readable watermarks in synthetic data can provide a clear signal for downstream models or validation tools. When a watermark is detected, the model or system could raise a flag or automatically classify the input as synthetic. However, this requires a recognized standard protocol for all generative models. 3) Fine-tuning with labeled synthetic Data. Including explicitly labeled synthetic and real examples during training (or fine-tuning) may help models learn the nuanced distinctions. However, this is inapplicable as it largely increases the cost and it is almost impossible to collect all generative models for training.
@@
Ethical issues of LVLMs. As existing LVLMs raise concerns about patient safety, liability, and public trust in AI-assisted medicine, we propose possible safeguards and regulatory considerations. Regulatory Oversight and Validation. Clinical Trials and Approval Processes: Before integrating LVLMs into clinical workflows, conducting rigorous validation studies—akin to clinical trials—can help quantify safety, efficacy, and error rates. Fail-Safe Mechanisms and Refusal to Answer. Confidence Estimation: Beyond the refusal-to-answer strategy seen in chatgpt, providing models with a mechanism to estimate their uncertainty can help them identify situations where they lack sufficient knowledge or clarity. This can be achieved through probabilistic outputs, confidence scores, or specialized OOD (Out-of-Distribution) detectors that signal when inputs deviate from the training domain. Human-in-the-Loop Verification: When the model’s confidence falls below a certain threshold, results should be flagged for clinician review. Ethical and Policy Recommendations. The consortium should establish accountability and liability for LVLMs. Clear guidelines are needed to define who is responsible if an LVLM’s recommendation leads to adverse patient outcomes.
Future perspectives and research direction. In addition to the current support for MRI data, our future endeavors aim to extend RadVUQA to include more imaging modalities—particularly ultrasound and X-ray. By incorporating these modalities, we hope to capture a broader spectrum of clinical scenarios and address modality-specific nuances, ultimately enhancing the generalizability and robustness of our benchmark. We plan to establish high-quality, annotated datasets for both ultrasound and X-ray, refining our evaluation methodologies to accommodate each modality’s distinct imaging characteristics. Additionally, longitudinal studies may be integrated into future iterations of RadVUQA, offering insights into disease progression and treatment responses over time. Beyond expanding modality coverage, interdisciplinary collaborations between AI experts and medical professionals will be crucial for developing clinically viable LVLMs. Close partnerships can help define clinically relevant tasks, annotate complex cases, and align model outputs with routine diagnostic workflows. Such synergy ensures that benchmarks like RadVUQA remain meaningful to end users, rather than purely academic exercises.
Meanwhile, emerging technologies such as explainable AI (XAI) [28, 29] and federated learning [30] hold significant promise for advancing LVLMs in real-world environments. Explainable AI offers pathways to interpret and visualize model decisions, which is essential for earning clinician trust and meeting regulatory expectations regarding transparency and accountability. Federated learning, on the other hand, addresses data privacy concerns by training models on distributed healthcare data without transferring sensitive patient information to a central server. This could both increase the diversity and volume of training data (boosting model robustness) while adhering to strict privacy regulations.
@@ VII Conclusion
  • LVLMs are not robust to unharmonized data, even certain models have lower hallucinations on the low-quality data.